OpenStudy (anonymous):

Write the equation of the line which passes through (5, –2) and is parallel to x = 4.

OpenStudy (anonymous):

y-(-2)=4(x-5) y+2=4x-20 y-4x=-22 4x-y=22

OpenStudy (anonymous):

equation of a line passing through point (x1,y1) and slope m is y-y1=m(x-x1) here m=4 (x1,y1)=(5,-2)

OpenStudy (anonymous):

fine?

OpenStudy (mayankdevnani):

@nitz WARNING=don't give direct answers....plz teach ...

OpenStudy (anonymous):

thank you!